OpenStudy (anonymous):

I need to know what the ordered pair is from 2x-6y=36 3x-9y=-27

jhonyy9 (jhonyy9):

2x-6y=36 multiply by -3 3x-9y=-27 multiply by 2 -6x+18y= - 108 6x-18y= -54 ---------------- 0 0 this sistem havent solutions